OpenStudy (anonymous):

A box with an open top and a square base is to hve a volume of 108 cubic inches. What dimensions allow for the least amount of material o construct the box? Calculus

OpenStudy (anonymous):

base's side: x height: h Volume: hx^2 = 108 h = 108/x^2 Surface area: 4hx + x^2 = x^2 + 432/x

OpenStudy (anonymous):

tell me if you don't understand anything surface area f(x) = x^2 + 432/x to find the minimum of this, we need to take the first derivative f'(x) = 2x - 432/x^2 then set f'(x) = 0 2x - 432/x^2 = 0 2x^3 - 432 = 0 x^3 = 216 x = 6 h = 108/x^2 = 108/36 = 3

OpenStudy (anonymous):

got it @npowers8 ?

OpenStudy (anonymous):

yes I am pretty sure. so the dimensions are 6 inches for base and 3 inches high

OpenStudy (anonymous):

yes